# From MRI to SARI: Parametric survey of non-axisymmetric accretion disk instabilities

### v1

@@ Abstract
Accretion disks are highly unstable to magnetic instabilities driven by shear flow, where classically, the axisymmetric, weak-field Magneto-Rotational Instability (MRI) has received much attention through local WKB approximations. In contrast, discrete non-axisymmetric counterparts require a more involved analysis through a full global approach to deal with the influence of the nearby magnetohydrodynamic (MHD) continua. Recently, rigorous MHD spectroscopy identified a new type of an ultra-localised, non-axisymmetric instability in global disks with super-Alfvénic flow. These Super-Alfvénic Rotational Instabilities (SARIs) fill vast unstable regions in the complex eigenfrequency plane with (near-eigen)modes that corotate at the local Doppler velocity and are radially localised between Alfvénic resonances. Unlike discrete modes, they are utterly insensitive to the radial disk boundaries. In this work, we independently confirm the existence of these unprecedented modes using our novel spectral MHD code Legolas, reproducing and extending our earlier study with detailed eigenspectra and eigenfunctions. We calculate growth rates of SARIs and MRI in a variety of disk equilibria, highlighting the impact of field strength and orientation, and find correspondence with analytical predictions for thin, weakly magnetised disks. We show that non-axisymmetric modes can significantly extend instability regimes at high mode numbers, with maximal growth rates comparable to the MRI. Furthermore, we explicitly show a region filled with quasi-modes whose eigenfunctions are extremely localised in all directions. These modes must be ubiquitous in accretion disks, and play a role in local shearing box simulations. Finally, we revisit recent dispersion relations in the Appendix, highlighting their relation to our global framework.
@@ 1 Introduction
Accretion disks are generally assumed to be in a turbulent state, as observed accretion rates are higher than expected from angular momentum transport through molecular viscosity alone (Pringle, 1981; Balbus & Hawley, 1998). The origin of this turbulence and resulting enhanced viscosity has long been unclear: hydrodynamic turbulence is largely excluded in typical accretion disk flow profiles that satisfy the Rayleigh criterion for linear stability. Non-axisymmetric hydrodynamic instability pathways, both linear and non-linear, for differentially rotating tori have been identified early on (Papaloizou & Pringle, 1984; Blaes & Hawley, 1988), but these more global modes are generally not expected to lead to fully developed turbulence in disks.
Only some thirty years ago, a convincing linear route to turbulence in disks was discovered through the weak-field, axisymmetric Magneto-Rotational Instability (MRI) (Balbus & Hawley, 1991), revisiting earlier analysis by Velikhov (1959) and Chandrasekhar (1960). Even a seed vertical magnetic field was shown to dramatically alter stability by acting as a tether between fluid elements, transporting angular momentum outwards. Non-linear simulations confirmed how at unstable MRI wavelengths, the quickly developing field distortions lead to angular momentum interchanges (Hawley & Balbus, 1991). Hawley & Balbus (1992) extended these non-linear 2D simulations in a shearing sheet to address their long-term evolution and found magnetic energy growth and sustained angular momentum transport. The linearly unstable, quickly growing MHD eigenmodes induce the generation of turbulent eddies at multiple scales, whose interactions may explain enhanced effective viscosity through the stress tensor. In the meantime, MRI has become an accepted pathway to disk turbulence through linear MHD instability in many fully 3D non-linear MHD simulations, from fully global Newtonian (Mishra et al., 2020) to general relativistic settings (Begelman et al., 2022; Ripperda et al., 2022), from local 3D shearing boxes in ideal or resistive MHD (Lesaffre et al., 2009; Bai & Stone, 2013; Hirai et al., 2018) or even kinetic particle-in-cell settings (Bacchini et al., 2022, 2024).
@@
In contrast to the axisymmetric MRI, non-axisymmetric modes fundamentally require such a global spectroscopic analysis because of singularities in the governing equations produced by the overlap of the continuous MHD spectra for radially varying, rotating equilibrium states. These non-axisymmetric instabilities are of particular interest for dynamo activity, which is impossible under axisymmetric conditions (Moffatt, 1978), and because their coupling to the toroidal field is essential for self-sustained instability (Hawley et al., 1996; Kitchatinov & Rüdiger, 2010; Begelman et al., 2022). Furthermore, although global, they can become very localised with growth rates equally large as the axisymmetric MRI, and may still be unstable under conditions where the MRI is typically stabilised (Goedbloed & Keppens, 2022; Begelman et al., 2022), as is also shown in this work. Additionally, secondary parasitic non-axisymmetric instabilities developing on channel flow solutions in shearing boxes have growth rates that can even exceed that of the MRI (Goodman & Xu, 1994). Although the original MRI description was immediately followed by a local WKB description of the non-axisymmetric MRI featuring transient growth (Balbus & Hawley, 1992), a systematic MHD spectroscopic approach has shown the existence of unstable modes that cannot be found from local theory alone (Matsumoto & Tajima, 1995; Ogilvie & Pringle, 1996; Goedbloed & Keppens, 2022). Additionally, growth rates for these ‘local’ non-axisymmetric MRI are one order of magnitude underestimated compared to their ‘global’ versions (Curry & Pudritz, 1996).
@@
Research on the global linear treatment of non-axisymmetric modes peaked in the decade after the MRI was discovered: Matsumoto & Tajima (1995) performed a global analysis of a local shearing sheet and found evanescent modes localised between two Alfvén resonances; Curry & Pudritz (1996) investigated a cylindrical equilibrium in the Boussinesq approximation with a constant vertical field and found two corotating Alfvénic modes, generalising the hydrodynamic Papaloizou-Pringle instability (Papaloizou & Pringle, 1984); Terquem & Papaloizou (1996) performed linear analysis and simulations of a cylindrical disk with both radial and vertical structure and limited to a toroidal field, and in their simulations obtained unstable modes growing on increasingly smaller scales; Ogilvie & Pringle (1996) compared non-axisymmetric modes in both a cylinder and shearing sheet for a purely toroidal field, and found two branches of modes clustering towards the continua in the limit $k\rightarrow+\infty$ with arbitrarily localised eigenfunctions; Noguchi et al. (2000) further explored the evanescent eigenmodes found by Matsumoto & Tajima (1995), although the range of unstable fields in their shearing sheet is at odds with the earlier cylindrical models; Keppens et al. (2002) gave a fully general theoretical and numerical description of the MHD eigenspectrum of a cylindrical model and found an abundance of unstable non-axisymmetric modes. This general formalism has yet to be applied to perform a comprehensive linear study of non-axisymmetric accretion disk instabilities, which is what we aim to accomplish in this work. In contrast to many of the above works, we will consider arbitrary magnetic field orientation and strength to show how the predictions from these limiting cases (purely vertical and azimuthal fields) hold up in general.
@@
In recent years, non-axisymmetric mode analysis has been sporadically revived, mostly in local shearing box models where the radial wavenumber changes in time and the instability is transient (Kitchatinov & Rüdiger, 2010; Mamatsashvili et al., 2013) or in WKB and global treatments of cylindrical models with a dominant azimuthal field (Das et al., 2018; Begelman et al., 2022) as it becomes clear that these modes must play an important role in the dynamics of accretion disks. However, in non-linear simulations, the development of turbulence is still often attributed solely to the axisymmetric MRI. Concurrently, there is claimed laboratory evidence for non-axisymmetric MRI in an azimuthal field (Seilmayer et al., 2014), for the standard vertical-field MRI (Wang et al., 2022b), and for unidentified non-axisymmetric modes (Wang et al., 2022a), to which linear global analysis was also applied (Ebrahimi & Pharr, 2022). All this proves that these are still exciting times for linear studies of non-axisymmetric instabilities, contrary to popular belief.
@@
Indeed, Goedbloed & Keppens (2022) recently found a completely new type of linear ideal MHD instability in thin, weakly magnetised, cylindrical accretion disks that could provide a broadly accessible pathway to turbulence. These modes densely populate a 2D region in the complex eigenfrequency plane and have eigenfunctions that resemble radially localised wave packages. These modes are promising because of the apparent ease with which they can be excited (they are not normal modes but require only a tiny addition of energy to be brought into resonance), their insensitivity to boundary conditions (the eigenfunctions are radially bound by Alfvén resonances) and the property that they exist all over the disk (each perturbation peaking at the local Doppler corotation radius). SARIs exist because of a delicate interplay between the MHD continua and the (near-)singularities they introduce in the analysis of the governing ordinary differential equations, hence they can only be obtained through a global analysis. Goedbloed & Keppens (2022) provided full analytic as well as numerical evidence for SARIs, using the powerful Spectral Web approach (Goedbloed, 2018a, b) to locate complex eigenfrequencies of ideal, stationary MHD states. In addition, they rigourously generalised a criterion for instability and growth rate prediction. In this work, we aim to confirm the existence of SARIs with an independent numerical approach and to show the applicability of this instability criterion.
@@
To distinguish between the transient growth of local non-axisymmetric MRI as described in the literature (Balbus & Hawley, 1992; Kitchatinov & Rüdiger, 2010; Mamatsashvili et al., 2013) and the global eigenmode nature of the non-axisymmetric SARIs, we will in the remainder of this work refer to any axisymmetric mode as the MRI, and to non-axisymmetric modes as SARIs in the regime of thin, weakly magnetised disks. We address some properties of non-axisymmetric modes that are known from local theory or limiting cases in global analysis but underexposed in the general global approach. In particular, we confirm analytical predictions for instability over a range of equilibria and show that the Alfvén frequency plays a central role, which opens up more regimes to instability. Furthermore, we show that global high-$m$ modes are nevertheless very localised, with possible implications for shearing box simulations.
@@
In this work, we first revisit, augment and further validate the results of Goedbloed & Keppens (2022) using the Legolas code (Claes et al., 2020). The equilibrium set-up and numerical approach are described in Sect. 2. Section 3 summarises the properties of the novel quasi-continuum modes, as well as those of the standard MRI and discrete SARI modes in a global approach. It will become clear that resonances with the Alfvén continua play a crucial role for non-axisymmetric modes. Second, we perform a comprehensive study of the most unstable MRI and SARI growth rates in a range of equilibria. Sections 5.15.3 compare growth rates for the MRI and SARIs and show that non-axisymmetric modes can significantly extend regimes of instability to stronger fields. These results are shown to match with classical and new analytical predictions in Sect. 5.4, and a comparison with local dispersion relation solutions highlights the necessity of a global approach. Finally, we again focus on the full MHD eigenspectra in Sect. 5.5 to explicitly show a quasi-continuum at high azimuthal wavenumber, featuring an abundance of modes with significant growth rates that are global in nature but radially and vertically extremely localised. As additional material, Appendix A summarises the mathematical tools required for a global and local analysis. Appendix B compares the general local dispersion relation of Blokland et al. (2005) and other recent dispersion relations that have been used for non-axisymmetric modes (Das et al., 2018; Ebrahimi & Pharr, 2022).
@@ 2.1 Model equilibrium
Our basic model representing a cylindrical accretion disk is an ideal MHD equilibrium (Keppens et al., 2002; Blokland et al., 2005; Goedbloed & Keppens, 2022) that radially varies according to self-similar profiles (Spruit et al., 1987) and that is vertically invariant. The disk is modelled as an annulus with radius $r\in[r_{1},r_{2}]$, with a line source approximating the central gravitating object, $\mathbf{g}(\mathbf{r})\approx\frac{GM_{*}}{r^{2}}\mathbf{\hat{r}}$. The assumption of vertical invariance in the disk is valid if small vertical wavelengths (or large wavenumbers $k$) are considered, much smaller than the scale height $H=c_{S}/(\Omega_{1}r_{1})\ll r_{1}$ in a thin, Keplerian disk (Shakura & Sunyaev, 1973), where $c_{S}=\sqrt{\gamma p/\rho}$ is the sound speed and $\Omega_{1}$ the disk rotation frequency. From hereon, we eliminate all proper units from the variables through normalising by equilibrium values at the inner boundary $r_{1}$, all denoted by a subscript $1$. The normalisation is fixed by choosing new units for density $\rho_{1}$, distance $r_{1}$ and velocity $v_{K1}=\sqrt{{GM_{*}}/{r_{1}}}$ (the inner Keplerian rotation velocity). The non-dimensionalised radial equilibrium profiles for density, pressure, velocity and magnetic field are then given by,
@@
In terms of these free parameters, which fully determine an equilibrium, the constants at the inner boundary become,
@@
where the rotation frequency is obtained from the radial force balance equation. Note that, although the rotation power-law is Keplerian, the actual rotation frequency $\Omega_{1}$ is in general not Keplerian because of pressure and tension forces. For convenience, we denote $\varphi=\arctan{B_{\theta 1}/B_{z1}}$ for the angle between the equilibrium magnetic field and the $z$-axis. Varying the field orientation from poloidal (vertical) to toroidal (azimuthal) is hence achieved through varying $\varphi$ from $0^{\circ}$ to $90^{\circ}$.

### v2

# Parametric Survey of Nonaxisymmetric Accretion Disk Instabilities: Magnetorotational Instability to Super-Alfvénic Rotational Instability

@@ Abstract
Accretion disks are highly unstable to magnetic instabilities driven by shear flow, where classically, the axisymmetric, weak-field Magneto-Rotational Instability (MRI) has received much attention through local WKB approximations. In contrast, discrete non-axisymmetric counterparts require a more involved analysis through a full global approach to deal with the influence of the nearby magnetohydrodynamic (MHD) continua. Recently, rigorous MHD spectroscopy identified a new type of an ultra-localised, non-axisymmetric instability in global disks with super-Alfvénic flow. These Super-Alfvénic Rotational Instabilities (SARIs) fill vast unstable regions in the complex eigenfrequency plane with (near-eigen)modes that corotate at the local Doppler velocity and are radially localised between Alfvénic resonances. Unlike discrete modes, they are utterly insensitive to the radial disk boundaries. In this work, we independently confirm the existence of these unprecedented modes using our novel spectral MHD code Legolas, reproducing and extending our earlier study with detailed eigenspectra and eigenfunctions. We calculate growth rates of SARIs and MRI in a variety of disk equilibria, highlighting the impact of field strength and orientation, and find correspondence with analytical predictions for thin, weakly magnetised disks. We show that non-axisymmetric modes can significantly extend instability regimes at high mode numbers, with maximal growth rates comparable to those of the MRI. Furthermore, we explicitly show a region filled with quasi-modes whose eigenfunctions are extremely localised in all directions. These modes must be ubiquitous in accretion disks, and play a role in local shearing box simulations. Finally, we revisit recent dispersion relations in the Appendix, highlighting their relation to our global framework.
@@ 1 Introduction
Accretion disks are generally assumed to be in a turbulent state, as observed accretion rates are higher than expected from angular momentum transport through molecular viscosity alone (Pringle, 1981; Balbus & Hawley, 1998). The origin of this turbulence and the resulting enhanced viscosity has long been unclear: hydrodynamic turbulence is largely excluded in typical accretion disk flow profiles that satisfy the Rayleigh criterion for linear stability. Non-axisymmetric hydrodynamic instability pathways, both linear and non-linear, for differentially rotating tori were identified early on (Papaloizou & Pringle, 1984; Blaes & Hawley, 1988), but these more global modes are generally not expected to lead to fully developed turbulence in disks.
Only some thirty years ago, a convincing linear route to turbulence in disks was discovered through the weak-field, axisymmetric Magneto-Rotational Instability (MRI) (Balbus & Hawley, 1991), revisiting earlier analysis by Velikhov (1959) and Chandrasekhar (1960). Even a seed vertical magnetic field was shown to dramatically alter stability by acting as a tether between fluid elements, transporting angular momentum outward. Non-linear simulations confirmed how, at unstable MRI wavelengths, the quickly developing field distortions lead to angular momentum interchanges (Hawley & Balbus, 1991). Hawley & Balbus (1992) extended these non-linear 2D simulations in a shearing sheet to address their long-term evolution and found magnetic energy growth and sustained angular momentum transport. The linearly unstable, quickly growing MHD eigenmodes induce the generation of turbulent eddies at multiple scales, whose interactions may explain enhanced effective viscosity through the stress tensor. In the meantime, MRI has become an accepted pathway to disk turbulence through linear MHD instability in many fully 3D non-linear MHD simulations, from fully global Newtonian (Mishra et al., 2020) and general relativistic settings (Begelman et al., 2022; Ripperda et al., 2022) to local 3D shearing boxes in ideal or resistive MHD (Lesaffre et al., 2009; Bai & Stone, 2013; Hirai et al., 2018) or even kinetic particle-in-cell settings (Bacchini et al., 2022, 2024).
@@
In contrast to the axisymmetric MRI, non-axisymmetric modes fundamentally require such a global spectroscopic analysis because of singularities in the governing equations produced by the overlap of the continuous MHD spectra for radially varying, rotating equilibrium states. These non-axisymmetric instabilities are of particular interest for dynamo activity, which is impossible under axisymmetric conditions (Moffatt, 1978), and because their coupling to the toroidal field is essential for self-sustained instability (Hawley et al., 1996; Kitchatinov & Rüdiger, 2010; Begelman et al., 2022). Furthermore, although global, they can become very localised with growth rates as high as those of the axisymmetric MRI, and may still be unstable under conditions where the MRI is typically stabilised (Goedbloed & Keppens, 2022; Begelman et al., 2022), as is also shown in this work. Additionally, secondary parasitic non-axisymmetric instabilities developing on channel flow solutions in shearing boxes have growth rates that can even exceed that of the MRI (Goodman & Xu, 1994). Although the original MRI description was immediately followed by a local WKB description of the non-axisymmetric MRI featuring transient growth (Balbus & Hawley, 1992), a systematic MHD spectroscopic approach has shown the existence of unstable modes that cannot be found from local theory alone (Matsumoto & Tajima, 1995; Ogilvie & Pringle, 1996; Goedbloed & Keppens, 2022). Additionally, growth rates for these ‘local’ non-axisymmetric MRI are an order of magnitude underestimated as compared to those of their ‘global’ versions (Curry & Pudritz, 1996).
@@
Research on the global linear treatment of non-axisymmetric modes peaked in the decade after the MRI was discovered: Matsumoto & Tajima (1995) performed a global analysis of a local shearing sheet and found evanescent modes localised between two Alfvén resonances; Curry & Pudritz (1996) investigated a cylindrical equilibrium in the Boussinesq approximation with a constant vertical field and found two corotating Alfvénic modes, generalising the hydrodynamic Papaloizou-Pringle instability (Papaloizou & Pringle, 1984); Terquem & Papaloizou (1996) performed linear analysis and simulations of a cylindrical disk having both radial and vertical structure and limited to a toroidal field, and in their simulations obtained unstable modes growing on increasingly smaller scales; Ogilvie & Pringle (1996) compared non-axisymmetric modes in both a cylinder and a shearing sheet for a purely toroidal field, and found two branches of modes clustering towards the continua in the limit $k\rightarrow+\infty$ with arbitrarily localised eigenfunctions; Noguchi et al. (2000) further explored the evanescent eigenmodes found by Matsumoto & Tajima (1995), although the range of unstable fields in their shearing sheet is at odds with the earlier cylindrical models; and Keppens et al. (2002) gave a fully general theoretical and numerical description of the MHD eigenspectrum of a cylindrical model and found an abundance of unstable non-axisymmetric modes. This general formalism has yet to be applied to perform a comprehensive linear study of non-axisymmetric accretion disk instabilities, which is what we aim to accomplish in this work. In contrast to many of the above works, we will consider an arbitrary magnetic field orientation and strength to show how the predictions from these limiting cases (purely vertical and azimuthal fields) hold up in general.
@@
In recent years, non-axisymmetric mode analysis has been sporadically revived, mostly in local shearing box models where the radial wavenumber changes in time and the instability is transient (Kitchatinov & Rüdiger, 2010; Mamatsashvili et al., 2013) or in WKB and global treatments of cylindrical models with a dominant azimuthal field (Das et al., 2018; Begelman et al., 2022) as it becomes clear that these modes must play an important role in the dynamics of accretion disks. However, in non-linear simulations, the development of turbulence is still often attributed solely to the axisymmetric MRI. Concurrently, there is claimed laboratory evidence for non-axisymmetric MRI in an azimuthal field (Seilmayer et al., 2014), for the standard vertical-field MRI (Wang et al., 2022b), and for unidentified non-axisymmetric modes (Wang et al., 2022a), to which linear global analysis has also been applied (Ebrahimi & Pharr, 2022). All this proves that these are still exciting times for linear studies of non-axisymmetric instabilities, contrary to popular belief.
@@
Indeed, Goedbloed & Keppens (2022) recently found a completely new type of linear ideal MHD instability in thin, weakly magnetised, cylindrical accretion disks that could provide a broadly accessible pathway to turbulence. These modes densely populate a 2D region in the complex eigenfrequency plane and have eigenfunctions that resemble radially localised wave packages. These modes are promising because of the apparent ease with which they can be excited (they are not normal modes but require only a tiny addition of energy to be brought into resonance), their insensitivity to boundary conditions (the eigenfunctions are radially bound by Alfvén resonances) and the property that they exist all over the disk (each perturbation peaking at the local Doppler corotation radius). SARIs exist because of a delicate interplay between the MHD continua and the (near-)singularities they introduce in the analysis of the governing ordinary differential equations, hence they can only be obtained through a global analysis. Goedbloed & Keppens (2022) provided full analytic as well as numerical evidence for SARIs, using the powerful Spectral Web approach (Goedbloed, 2018a, b) to locate complex eigenfrequencies of ideal, stationary MHD states. In addition, they rigorously generalised a criterion for instability and growth rate prediction. In this work, we aim to confirm the existence of SARIs with an independent numerical approach and to show the applicability of this instability criterion.
@@
To distinguish between the transient growth of local non-axisymmetric MRI as described in the literature (Balbus & Hawley, 1992; Kitchatinov & Rüdiger, 2010; Mamatsashvili et al., 2013) and the global eigenmode nature of the non-axisymmetric SARIs, we will in the remainder of this work refer to any axisymmetric mode as the MRI, and to non-axisymmetric modes as SARIs in the regime of thin, weakly magnetised disks. We address some properties of non-axisymmetric modes that are known from local theory or limiting cases in global analysis but are underexposed in the general global approach. In particular, we confirm analytical predictions for instability over a range of equilibria and show that the Alfvén frequency plays a central role, which opens up more regimes to instability. Furthermore, we show that global high-$m$ modes are nevertheless very localised, with possible implications for shearing box simulations.
@@
In this work, we first revisit, augment, and further validate the results of Goedbloed & Keppens (2022) using the Legolas code (Claes et al., 2020). The equilibrium set-up and numerical approach are described in Sect. 2. Section 3 summarises the properties of the novel quasi-continuum modes, as well as those of the standard MRI and discrete SARI modes in a global approach. It will become clear that resonances with the Alfvén continua play a crucial role in the behaviour of non-axisymmetric modes. Second, we perform a comprehensive study of the most unstable MRI and SARI growth rates in a range of equilibria. Sections 5.15.3 compare growth rates for the MRI and SARIs and show that non-axisymmetric modes can significantly extend regimes of instability to stronger fields. These results are shown to match classical and new analytical predictions in Sect. 5.4, and a comparison with local dispersion relation solutions highlights the necessity of a global approach. Finally, we again focus on the full MHD eigenspectra in Sect. 5.5 to explicitly show a quasi-continuum at a high azimuthal wavenumber, featuring an abundance of modes with significant growth rates that are global in nature but radially and vertically extremely localised. As additional material, Appendix A summarises the mathematical tools required for a global and local analysis. Appendix B compares the general local dispersion relation of Blokland et al. (2005) and other recent dispersion relations that have been used for non-axisymmetric modes (Das et al., 2018; Ebrahimi & Pharr, 2022).
@@ 2.1 Model equilibrium
Our basic model representing a cylindrical accretion disk is an ideal MHD equilibrium (Keppens et al., 2002; Blokland et al., 2005; Goedbloed & Keppens, 2022) that radially varies according to self-similar profiles (Spruit et al., 1987) and that is vertically invariant. The disk is modelled as an annulus with radius $r\in[r_{1},r_{2}]$, with a line source approximating the central gravitating object, $\mathbf{g}(\mathbf{r})\approx\frac{GM_{*}}{r^{2}}\mathbf{\hat{r}}$. The assumption of vertical invariance in the disk is valid if small vertical wavelengths (or large wavenumbers $k$) are considered, much smaller than the scale height $H=c_{S}/(\Omega_{1}r_{1})\ll r_{1}$ in a thin, Keplerian disk (Shakura & Sunyaev, 1973), where $c_{S}=\sqrt{\gamma p/\rho}$ is the sound speed and $\Omega_{1}$ the disk rotation frequency. From here on, we eliminate all proper units from the variables through the equilibrium values at the inner boundary $r_{1}$, all denoted by a subscript $1$. The normalisation is fixed by choosing new units for density $\rho_{1}$, distance $r_{1}$ and velocity $v_{K1}=\sqrt{{GM_{*}}/{r_{1}}}$ (the inner Keplerian rotation velocity). The non-dimensionalised radial equilibrium profiles for the density, pressure, velocity, and magnetic field are then given by
@@
In terms of these free parameters, which fully determine an equilibrium, the constants at the inner boundary become
@@
where the rotation frequency is obtained from the radial force balance equation. Note that, although the rotation power-law is Keplerian, the actual rotation frequency $\Omega_{1}$ is in general not Keplerian because of pressure and tension forces. For convenience, we denote $\varphi=\arctan{B_{\theta 1}/B_{z1}}$ for the angle between the equilibrium magnetic field and the $z$-axis. Variation of the field orientation from poloidal (vertical) to toroidal (azimuthal) is hence achieved through varying $\varphi$ from $0^{\circ}$ to $90^{\circ}$.